Please help.
Find all the values of x satisfying the equation.
6+ 1/x+1 = 2/x+2
x=__
If there is more than one solution, separate them with commas.

Answer by edjones(8007)

6+ 1/(x+1) = 2/(x+2) I assume that this is the problem.
6(x+1)(x+2) + (x+2) = 2(x+1)
6x^2+18x+12+x+2=2x+2
6x^2+18x+12+x+2-2x-2=0
6x^2+17x+12=0
6*12=72 what two factors of 72 have a sum of 17? 8 & 9
6x^2+9x+8x+12=0
3x(2x+3)+4(2x+3)=0
(3x+4)(2x+3)=0
.
3x=-4
x=-4/3
.
2x=-3
x=-3/2
.
Ed